# A Survey of Quantized Graph Representation Learning: Connecting Graph Structures with Large Language Models

### v1

@@ Abstract
Recent years have witnessed rapid advances in graph representation learning, with the continuous embedding approach emerging as the dominant paradigm. However, such methods encounter issues regarding parameter efficiency, interpretability, and robustness. Thus, Quantized Graph Representation (QGR) learning has recently gained increasing interest, which represents the graph structure with discrete codes instead of conventional continuous embeddings. Given its analogous representation form to natural language, QGR also possesses the capability to seamlessly integrate graph structures with large language models (LLMs). As this emerging paradigm is still in its infancy yet holds significant promise, we undertake this thorough survey to promote its rapid future prosperity. We first present the background of the general quantization methods and their merits. Moreover, we provide an in-depth demonstration of current QGR studies from the perspectives of quantized strategies, training objectives, distinctive designs, knowledge graph quantization, and applications. We further explore the strategies for code dependence learning and integration with LLMs. At last, we give discussions and conclude future directions, aiming to provide a comprehensive picture of QGR and inspire future research.
@@ 1 Introduction
Graph representation usually involves transforming nodes, edges, or structures within graph data to low-dimensional dense embeddings, which is called continuous graph representation as shown in Figure 1 (a). The learned representation is obligated to preserve both node attributes and topological structure Zhu et al. (2020). This field has seen substantial development and expansion in recent years. The emergence of numerous advanced technologies has invariably drawn significant attention within the graph community, becoming focal points of extensive studies, e.g., the well-known node2vec Grover and Leskovec (2016) model, graph neural networks (GNNs) Kipf and Welling (2017), and self-supervised graph learning Liu et al. (2022). They have achieved great empirical success in graph-related domains, including bioinformatics, social networks, recommendation systems, and anomaly detection Xia et al. (2021); Liu et al. (2023c).
@@
Although continuous graph representation successfully extracts rich semantic information from the input graph to adapt to a wide range of machine learning and deep learning tasks, concerns may arise regarding the embedding efficiency, interpretability, and robustness Wang et al. (2024). Specifically, with the great growth of graph scales, there will be a linear increase in embedding parameters, becoming a considerable challenge in graph learning contexts involving millions or even more nodes. Besides, continuous embeddings are generated in a black-box fashion and the meaning of the overall representation or any one dimension is unknowable, leading to a lack of interpretability. Moreover, these continuous representations are typically task-specific and show no significant abstraction at the representation level, making them not robust enough to generalize to a variety of practical tasks as effectively as LLMs.
@@
Interestingly, in real life, discrete features are often preferred for representation. For example, as shown in Figure 1 (b), descriptions affiliated to an individual are often a combination of discrete and meaningful features, such as age (e.g., 18), gender (male or female), and place of residence (e.g., Beijing). This scheme is simple and easy to understand. Based on this intuition, quantized representation came into being, which is now a hot topic in the computer vision domain that can facilitate image generation by combining with Transfomer or LLMs Esser et al. (2021); Lee et al. (2024). Inspired by general quantized representation techniques, quantized graph representation (QGR) learning has recently been proposed to address the aforementioned shortcomings of continuous graph learning. The core idea is to learn discrete codes for nodes, subgraphs, or the overall graph structures, instead of employing continuous embeddings. As shown in Figure 1 (b), a node can be represented as discrete codes of (male, 18, Beijing, Ph.D.). In QGR methods, codes can be learned by differentiable optimization with specific graph targets or self-supervised graph learning. It has the capacity to learn more high-level representations due to a compact latent space, which would enhance the representation efficiency, interpretability, and robustness. Recall that when faced with graph-text scenarios, incorporating continuous embeddings into advanced LLMs is inherently challenging and prone to introducing the loss of semantic information. In contrast, due to the consistency with the discrete nature of natural language, the output of QGR would be seamlessly integrated into LLMs, like early-fusion strategy Team (2024). Considering that this emerging paradigm is still in its initial stage but carries substantial potential in the era of LLMs, we undertake this thorough survey. It encompasses various viewpoints, including different graph types, quantized strategies, training objectives, distinctive designs, applications and so forth. We also provide discussions and conclude future directions, intending to motivate future investigation and contribute to the advancement of the graph community.
@@
To start with, we first briefly give the background in Section 2. Then, we provide a comprehensive introduction to QGR frameworks from multiple perspectives in Section 3. Next, we present the code dependence learning as well as the integration with LLMs in Section 4 and Section 5, respectively. Finally, we give the discussions and future directions in Section 6.
2 Background
2.1 Merits of Quantized Graph Representation
Embedding Parameter Efficiency.
As large-scale graphs become more prevalent, there is a corresponding substantial increase in the representation parameters, which necessitates significant memory usage for storage and substantial computational expenses Galkin et al. (2022). Supposing as the number of nodes and as the dimensions, the required embedding parameters would be . In the quantized settings, the total count of parameters would reduce to , where and denote the length of the codeword set and the number of codes assigned to each node. For example, if there are nodes with 1024 feature dimensions, 1024 codewords, and 8 assigned codes, it would be a 113-fold reduction in required embedding parameters.
Explainability and Interpretability.
The discrete nature of QGR brings the explainability and interpretability for embeddings and reasoning, as shown in Figure1 (b). Each code would be assigned to the practical item by specific designs for direct interpretability, e.g., Dr. E Liu et al. (2024b) introduces a language vocabulary as the code set and each node in graphs can be represented as the permutation of explainable language tokens.
Robustness and Generalization.
The discrete tokens of the QGR are more robust and generalizable, which usually utilizes the self-supervised strategies to learn both local-level and high-level graph structures (e.g., edge reconstruction and DGI Velickovic et al. (2019)) as well as semantics (e.g., feature reconstruction). Moreover, the acquired tokens can readily adapt to downstream applications, either directly or indirectly Yang et al. (2024).
Seamless Integration with NLP Models.
The swift advancement in natural language processing (NLP) techniques, particularly LLMs, has sparked an increased interest in employing NLP models to resolve graph tasks. However, the inherent representation gap between the typical graph structure and natural language poses a significant challenge to their seamless and effective integration. Graph quantization, by learning discrete codes that bear resemblance to natural language forms, can facilitate this integration directly and seamlessly Lin et al. (2025a); Liu et al. (2024b).
2.2 General Quantization Methods
@@
The mainstream quantization methods can be roughly categorized into product quantization, vector quantization, finite scalar quantization, and anchor selection and assignment.
Product quantization (PQ).
@@
The core idea of PQ Jegou et al. (2011) is to divide a high-dimensional space into multiple low-dimensional subspaces and then quantize them independently within each subspace. Specifically, it splits a high-dimensional vector into multiple smaller subvectors and quantizes each subvector separately. In the quantization process, each subvector is mapped to a set of a finite number of center points by minimizing the quantization error.
Vector Quantization (VQ).
@@
To effectively learn quantized representations in a differentiable manner, the VQ strategy van den Oord et al. (2017); Esser et al. (2021) is proposed and has emerged as the predominant method in the field. The core idea involves assigning the learned continuous representations to the codebook’s index and employing the Straight-Through Estimator Bengio et al. (2013) for effective optimization. The formulation is as follows. , , denote the encoder, decoder, and quantization process, respectively. is the codebook representation with dimension , which corresponds to discrete codewords . is the graph, where is the node set with size and is the edge set. The presence of or is not guaranteed, where is the input feature of a graph and is the label set of each node. is the latent representation after encoder and is the reconstruction feature after quantization and decoder, i.e., , .
@@
Using VQ techniques, a continuous vector can be quantized to the counterpart of a discrete code. However, there would be distinctions and differences in the vector pre and post-quantization. To address it, Residual Vector Quantization (RVQ) Lee et al. (2022) is proposed to constantly fit residuals caused by the quantization process. The whole process can be seen in Algorithm 1. RVQ is actually a multi-stage quantization method and usually has multiple codebooks. The codebook stores the residual values of the quantized vector and the original vector at each step, and the original vector can be approximated infinitely by quantizing the residual values at multiple steps.
Finite Scalar Quantization (FSQ).
Because of the introduction of the codebook, there would be complex technique designs and codebook collapse in VQ techniques. Thus, FSQ Mentzer et al. (2024) introduces a very simple quantized strategy, i.e., directly rounding to integers rather than explicitly introducing parameters of the codebook. FSQ projects the hidden representations into a few dimensions (usually fewer than 10) and each dimension is quantized into a limited selection of predetermined values, which inherently formulates a codebook as a result of these set combinations. It is formulated as:
@@
Anchor Selection and Assignment (ASA).
It is typically unsupervised and often employs prior strategies to identify informative nodes within the graph, e.g., the Personalized PageRank Page (1999) and node degree. Consequently, each node can be attributed to a combination of anchors that exhibit structural or semantic relationships, like using the shortest path. This form of quantization designates the code as a real node, unlike the aforementioned three methods.
@@ 3 Quantized Graph Learning Framework
The general process of QGR is illustrated in Figure 2, which comprises the encoder , decoder , and quantization process . The encoder is to model the original graph structure into latent space, where MLPs He et al. (2020) and GNNs Xia et al. (2023) are usually utilized. The decoder also commonly utilizes GNNs for structure embedding and the quantization process implements the strategies outlined in Section 2.2. In particular, Bio2Token Liu et al. (2024a) utilizes Mamba Gu and Dao (2023) as both encoder and decoder. UniMoT Zhang et al. (2024) leverages the pre-trained MoleculeSTM molecule encoder Liu et al. (2023a) to connect the molecule graph encoder and causal Q-Former Li et al. (2023b). In the following sections, we will delve into the details of the model design.
Model | one/p | Self-Sup | Graph Type | Quantization | Training | Application | Dependence | w/ LM |
[…]
@@ 3.1 Quantized Strategies
Except for SNEQ He et al. (2020) and d-SNEQ He et al. (2023), the majority of research tends to employ VQ-related strategies. Beyond the general techniques in VQ, Dr.E Liu et al. (2024b) utilizes RVQ in one specific layer, which is called intra-layer residue. For more effective in encoding the structural information of the central node, it involves preserving the multi-view of a graph. Specifically, between GCN layers, the inter-layer residue is carried out to enhance the representations:
where and denote the latent representations and quantized ones. is the number of learned codes for each layer. Using this process, the discrete codes of each node are generated in a gradual (auto-regressive) manner, following two specific orders: from previous to subsequent, and from the bottom layer to the top. This approach guarantees comprehensive multi-scale graph structure modeling and efficient code learning. iMoLD Zhuang et al. (2023) proposes another type of Redidual VQ, incorporating both the continuous and discrete representations, i.e., , as the final node representation.
GLAD Boget et al. (2024b) and Bio2Token Liu et al. (2024a) utilize the straightforward approach, FSQ, to facilitate the learning of discrete codes. For KG quantization, early methods such as NodePiece Galkin et al. (2022), EARL Chen et al. (2023), and RandomEQ Li et al. (2023a) have adopted the unsupervised learning approach ASA to achieve effective learning outcomes. For more condensed code representations, SSQR Lin et al. (2025a) employs the VQ strategy to capture both the structures and semantics of KGs.
@@ 3.2 Training Objectives
Node Level.
Feature reconstruction usually serves as a simple self-supervised method for QGR learning, which is to compare the original node features with reconstructed features based on the learned codes. It mainly has the following two implementation types:
@@
Edge Level.
It is a primary strategy to acquire the graph structures in QGR codes using a self-supervised paradigm. The direct link prediction, i.e., edge reconstruction, is a widely adopted technique to assess the presence or absence of each edge before and after the reconstruction process as:
@@
Graph Level.
For high-level representations, the graph modeling targets are proposed. For instance, graph contrastive learning is a widely utilized method for modeling relationships between graph pairs:
@@ 3.3 Distinctive Designs
Codebook Design.
Beyond randomly setting the codebook for updating, there are several specifically designed strategies for domain knowledge adaption and generalization. For example, Dr.E Liu et al. (2024b) sets the vocabulary of LLaMA as the codebook, realizing token-level alignment between GNNs and LLM. In this way, each node of graphs can be quantized to a permutation of language tokens and then can be directly input to LLMs to make predictions. When modeling molecular graphs, Mole-BERT Xia et al. (2023) categorizes the codebook embeddings into various groups, each representing a distinct type of atom. For instance, the quantized codes of carbon, nitrogen and oxygen are confined to different groups.
Training Pipelines.
Aside from the single-stage learning process applied to QGR and particular tasks, some methods incorporate multiple stages into the training framework. GQT Wang et al. (2024) modulate the learned codes through hierarchical encoding and structural gating, which are subsequently fed into the Transformer network and aggregate the learned representations through an attention module. Based on the learned quantized representation, NID Luo et al. (2024) trains an MLP network for downstream tasks, such as node classification and link prediction.

### v2

@@ Abstract
Recent years have witnessed rapid advances in graph representation learning, with the continuous embedding approach emerging as the dominant paradigm. However, such methods encounter issues regarding parameter efficiency, interpretability, and robustness. As a promising solution, Quantized Graph Representation (QGR) learning has recently gained increasing interest, which encodes graph structures with discrete codes instead of conventional continuous embeddings. Given its analogous representation form to natural language, QGR also possesses the capability to seamlessly integrate graph structures with large language models (LLMs). As this emerging paradigm is still in its infancy yet holds significant promise, we undertake this thorough survey to promote its rapid future prosperity. We first present the background of the general quantization methods and their merits. An in-depth demonstration of existing QGR studies is then carried out from the perspectives of quantized strategies, training objectives, distinctive designs, knowledge graph quantization, and applications. Furthermore, we explore the strategies for code dependence learning and integration with LLMs. The final discussion on strengths and limitations aims to provide a comprehensive picture of QGR and inspire future research directions.
@@ 1 Introduction
Graph representation usually involves converting nodes, edges, or structures within various graph data into low-dimensional dense embeddings, which is called continuous graph representation Xia et al. (2021); Lin et al. (2023) as shown in Figure 1 (a). The learned representation is obligated to preserve both node attributes and topological structures Zhu et al. (2020). This field has made great progress in recent years, and the emergence of numerous advanced technologies has invariably drawn significant attention within the graph community, becoming focal points of extensive studies, e.g., the well-known node2vec Grover and Leskovec (2016) model, graph neural networks (GNNs) Kipf and Welling (2017), and self-supervised graph learning Liu et al. (2022). They have achieved significant empirical success in graph-related domains, including social networks, recommendation systems, and anomaly detection Xia et al. (2021); Liu et al. (2023c).
@@
Although continuous graph representation successfully extracts rich semantic information from the input graph to adapt to a wide range of machine learning and deep learning tasks, concerns may arise regarding the embedding efficiency, interpretability, and robustness Wang et al. (2024). Specifically, with the great growth of graph scales, there will be a linear increase in embedding parameters, becoming a considerable challenge in graph learning contexts involving millions or even more nodes. Besides, continuous embeddings are generated in a black-box fashion and the meaning of the overall representation or any one dimension is unknowable, leading to a lack of interpretability. Moreover, these continuous representations are typically task-specific and show no significant abstraction at the representation level, making them not robust enough to generalize to a variety of practical tasks as effectively as large language models (LLMs) Naveed et al. (2025).
Interestingly, in real life, discrete features are often preferred for representation. For example, as shown in Figure 1 (b), descriptions affiliated to an individual are often a combination of discrete and meaningful features, such as age (e.g., 18), gender (male or female), and place of residence (e.g., Beijing). This scheme is simple and easy to understand. Based on this intuition, quantized representation came into being, which is now a hot topic in the computer vision domain that can facilitate image generation by combining with Transfomer or LLMs Esser et al. (2021); Lee et al. (2024).
@@
Inspired by general quantized representation techniques, quantized graph representation (QGR) learning has recently been proposed to address the aforementioned shortcomings of continuous graph learning. The core idea is to learn discrete codes for nodes, subgraphs, or the overall graph structures, instead of employing continuous embeddings. As shown in Figure 1 (b), a node can be represented as discrete codes of (male, 18, Beijing, Ph.D.). In QGR methods, codes can be learned by differentiable optimization with specific graph targets or self-supervised strategies, illustrated in Figure 2. It has the capacity to learn more high-level representations due to a compact latent space, which would enhance the representation efficiency, interpretability, and robustness. Recall that when faced with graph-text scenarios, incorporating continuous embeddings into advanced LLMs is inherently challenging and prone to introducing the loss of semantic information. In contrast, due to the consistency with the discrete nature of natural language, the output of QGR would be seamlessly integrated into LLMs, like early-fusion strategy Team (2024). Considering that this emerging paradigm is still in its initial stage but carries substantial potential in the era of LLMs, we undertake this thorough survey. It encompasses various viewpoints, including different graph types, quantized strategies, training objectives, distinctive designs, applications and so forth. We also conclude future directions, intending to motivate future investigation and contribute to the advancement of the graph and natural language community.
@@
To start with, we first describe the general quantization method in §2. Then, we provide a comprehensive introduction to QGR frameworks from multiple perspectives in §3. Next, we present the code dependence learning as well as the integration with LLMs in §4 and §5. Finally, we give future directions in §6.
2 General Quantization Methods
@@
We provide a detailed summary of the merits of QGR in Appendix §A. The mainstream quantization methods can be roughly categorized into product quantization, vector quantization, finite scalar quantization, and anchor selection and assignment.
@@
Product quantization (PQ). The core idea of PQ Jegou et al. (2011) is to divide a high-dimensional space into multiple low-dimensional subspaces and then quantize them independently within each subspace. Specifically, it splits a high-dimensional vector into multiple smaller subvectors and quantizes each subvector separately. In the quantization process, each subvector is mapped to a set of a finite number of center points by minimizing the quantization error.
@@
Vector Quantization (VQ). To effectively learn quantized representations in a differentiable manner, the VQ strategy van den Oord et al. (2017); Esser et al. (2021) is proposed and has emerged as the predominant method in the field. The core idea involves assigning the learned continuous representations to the codebook’s index and employing the Straight-Through Estimator Bengio et al. (2013) for effective optimization. The formulation is as follows. , , denote the encoder, decoder, and quantization process, respectively. is the codebook representation with dimension , which corresponds to discrete codewords . is the graph, where is the node set with size and is the edge set. The presence of or is not guaranteed, where is the input feature of a graph and is the label set of each node. is the latent representation after encoder and is the reconstruction feature after quantization and decoder, i.e., , .
@@
| Model | one/p | Self-Sup | Graph Type | Method | Training | Application | Dependence | w/ LM |
| SNEQ He et al. (2020) | ✓ | AG (HOG) | PQ | STC, SEC | NC, LP, NR | ✗ | ✗ | |
| d-SNEQ He et al. (2023) | ✓ | AG (HOG) | PQ | STC, SEC | NC, LP, NR, PP | ✗ | ✗ | |
| Mole-BERT Xia et al. (2023) | ✓ | MG (HOG) | VQ | FR/MCM, GCL | GC, GR | ✓ | ✗ | |
| iMoLD Zhuang et al. (2023) | ✓ | MG (HOG) | VQ | IL, GC (GR) | GC, GR | ✗ | ✗ | |
| VQGraph Yang et al. (2024) | ✓ | HOG | VQ | FR, LP/NC, Distill | NC | ✗ | ✗ | |
| Dr.E Liu et al. (2025) | ✓ | TAG (HOG) | RVQ | FR, LP, NC/FR, NTP | NC | ✓ | ✓ | |
| STAG Bo et al. (2025) | ✓ | TAG (HOG) | VQ | GraphMAE, Contrast/ | NC | ✗ | ✓ | |
| DGAE Boget et al. (2024) | ✓ | HEG | VQ | LP/NTP | GG | ✓ | ✗ | |
| LLPS Donà et al. (2025) | ✓ | MG (HOG) | FSQ | Reconstruction | GG | ✗ | ✗ | |
| GLAD Nguyen et al. (2025) | ✓ | MG (HOG) | FSQ | FR+LP/Diffusion Bridge | GG | ✓ | ✗ | |
| Bio2Token Liu et al. (2024) | ✓ | MG (HOG) | FSQ | FR | GG | ✗ | ✗ | |
| GQT Wang et al. (2024) | ✓ | HOG | RVQ | DGI, GraphMAE2/NC | NC | ✗ | ✗ | |
| NID Luo et al. (2024) | ✓ | HOG | RVQ | GraphMAE (GraphCL)/ | NC, LP, GC, GR, etc | ✗ | ✗ | |
| UniMoT Zhang et al. (2024) | ✓ | MG (HOG) | VQ | FR/NTP | GC, GR, Captioning, etc. | ✗ | ✓ | |
| NodePiece Galkin et al. (2022) | ✗ | KG | ASA | –/LP | LP | ✗ | ✗ | |
| EARL Chen et al. (2023) | ✗ | KG | ASA | –/LP | LP | ✗ | ✗ | |
| RandomEQ Li et al. (2023a) | ✗ | KG | ASA | –/LP | LP | ✗ | ✗ | |
| SSQR Lin et al. (2025a) | ✓ | KG | VQ | LP, Distilling/NTP | LP, Triple Classification | ✗ | ✓ | |
| ReaLM Guo et al. (2025) | ✓ | KG | RVQ | FR/NTP | LP, Triple Classification | ✗ | ✓ | |
| MedTok Su et al. (2025) | ✓ | KG | VQ | Contrast/ | Classification, QA | ✗ | ✓ |
Using VQ techniques, a continuous vector can be quantized to the counterpart of a discrete code. However, there would be distinctions and differences in the vector pre and post-quantization. To address it, Residual Vector Quantization (RVQ) Lee et al. (2022) is proposed to constantly fit residuals caused by the quantization process. RVQ is actually a multi-stage quantization method and usually has multiple codebooks. The codebook stores the residual values of the quantized vector and the original vector at each step, and the original vector can be approximated infinitely by quantizing the residual values at multiple steps. The whole process is shown in Algorithm 1 of the Appendix.
Finite Scalar Quantization (FSQ). Because of the introduction of the codebook, there would be complex technique designs and codebook collapse in VQ techniques. Thus, FSQ Mentzer et al. (2024) introduces a very simple quantized strategy, i.e., directly rounding to integers rather than explicitly introducing parameters of the codebook. FSQ projects the hidden representations into a few dimensions (usually fewer than 10) and each dimension is quantized into a limited selection of predetermined values, which inherently formulates a codebook as a result of these set combinations. It is formulated as:
@@
Anchor Selection and Assignment (ASA). It is typically unsupervised and often employs prior strategies to identify informative nodes within the graph, e.g., the Personalized PageRank Page (1999) and node degree. Consequently, each node can be attributed to a combination of anchors that exhibit structural or semantic relationships, like using the shortest path. This form of quantization designates the code as a real node, unlike the aforementioned three methods.
@@ 3 Quantized Graph Learning Framework
The general process of QGR is illustrated in Figure 2, which comprises the encoder , decoder , and quantization process . The encoder is to model the original graph structure into latent space, where MLPs He et al. (2020) and GNNs Xia et al. (2023) are usually utilized. The decoder also commonly utilizes GNNs for structure embedding and the quantization process implements the strategies outlined in Section 2. In particular, Bio2Token Liu et al. (2024) utilizes Mamba Gu and Dao (2023) as both encoder and decoder. UniMoT Zhang et al. (2024) leverages the pre-trained MoleculeSTM molecule encoder Liu et al. (2023a) to connect the molecule graph encoder and causal Q-Former Li et al. (2023b). In the following sections, we will delve into the details of the model design.
@@ 3.1 Quantized Strategies
Except for SNEQ He et al. (2020) and d-SNEQ He et al. (2023), the majority of research tends to employ VQ-related strategies. Beyond the general techniques in VQ, Dr.E Liu et al. (2025) utilizes RVQ in one specific layer, which is called intra-layer residue. For more effective in encoding the structural information of the central node, it involves preserving the multi-view of a graph. Specifically, between GCN layers, the inter-layer residue is carried out to enhance the representations:
where and denote the latent representations and quantized ones. is the number of learned codes for each layer. Using this process, the discrete codes of each node are generated in a gradual (auto-regressive) manner, following two specific orders: from previous to subsequent, and from the bottom layer to the top. This approach guarantees comprehensive multi-scale graph structure modeling and efficient code learning. iMoLD Zhuang et al. (2023) proposes another type of Residual VQ, incorporating both the continuous and discrete representations, i.e., , as the final node representation.
GLAD Nguyen et al. (2025) and Bio2Token Liu et al. (2024) utilize the straightforward approach, FSQ, to facilitate the learning of discrete codes. For KG quantization, early methods such as NodePiece Galkin et al. (2022), EARL Chen et al. (2023), and RandomEQ Li et al. (2023a) have adopted the unsupervised learning approach ASA to achieve effective learning outcomes. For more condensed code representations, SSQR Lin et al. (2025a) employs the VQ strategy to capture both the structures and semantics of KGs.
@@ 3.2 Training Objectives
Node Level. Feature reconstruction usually serves as a simple self-supervised method for QGR learning, which is to compare the original node features with reconstructed features based on the learned codes. It mainly has the following two implementation types:
@@
Edge Level. It is a primary strategy to acquire the graph structures in QGR codes using a self-supervised paradigm. The direct link prediction, i.e., edge reconstruction, is a widely adopted technique to assess the presence or absence of each edge before and after the reconstruction process as:
@@
Graph Level. For high-level representations, the graph modeling targets are proposed. For instance, graph contrastive learning van den Oord et al. (2018) is a widely utilized method for modeling relationships between graph pairs:
@@ 3.3 Distinctive Designs
Codebook Design. Beyond randomly setting the codebook for updating, there are several specifically designed strategies for domain knowledge adaption and generalization. For example, Dr.E Liu et al. (2025) sets the vocabulary of LLaMA as the codebook, realizing token-level alignment between GNNs and LLM. In this way, each node of graphs can be quantized to a permutation of language tokens and then can be directly input to LLMs to make predictions. When modeling molecular graphs, Mole-BERT Xia et al. (2023) categorizes the codebook embeddings into various groups, each representing a distinct type of atom. For instance, the quantized codes of carbon, nitrogen and oxygen are confined to different groups.
Training Pipelines. Aside from the single-stage learning process applied to QGR and particular tasks, some methods incorporate multiple stages into the training framework. GQT Wang et al. (2024) modulate the learned codes through hierarchical encoding and structural gating, which are subsequently fed into the Transformer network and aggregate the learned representations through an attention module. Based on the learned quantized representation, NID Luo et al. (2024) trains an MLP network for downstream tasks, such as node classification and link prediction.